A spherical ball of plasma has a radius of 0.50 m. The electric field on the surface of the ball is measured to be 8.9 x 103 N/C and points radially inward toward the center. What is the value of the electric charge


1
Expert's answer
2022-02-22T04:41:35-0500

The electric field

"E=k\frac{q}{r^2}"

Hence, the charge of a ball

"q=Er^2/k\\
=8.9*10^3*0.50^2/9*10^9=2.5*10^{-7}\:\rm C"